OpenStudy (anonymous):

Suppose a triangle has two sides of length 2 and 3 and that the angle between these two sides is pi/3. What is the length of the third side of the triangle? ...is the answer sqrt7?

OpenStudy (anonymous):

Use Law of Cosines: Let a,b,c be the side lengths and let C be the angle between a and b. Then c^2 = a^2 + b^2 - 2ab cos C Now plug in a = 2, b = 3, C = pi/3 c^2 = 2^2 + 3^2 - 2(2)(3) cos (pi/3) c^2 = 4 + 9 - 12*(1/2) c^2 = 4 + 9 - 6 c^2 = 7

OpenStudy (anonymous):

So c = sqrt(7)

OpenStudy (anonymous):

Yea @FutureMathProfessor Beat Me To It xD

OpenStudy (anonymous):

I'll still give you a medal though..as long as @Exvikki gives me a medal! xD

OpenStudy (anonymous):

Gave You One Right Back Lol

OpenStudy (anonymous):

<3 Thank you :D

OpenStudy (anonymous):

Any time babe <3